OpenStudy (anonymous):

how do you solve x/7 - 8=-7 x/7= x over 7

OpenStudy (anonymous):

x/7 - 8 = -7 x/7 = -7 + 8 x/7 = 1 x = 7

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

but it;s kind of confusing

OpenStudy (anonymous):

What is confusing?

OpenStudy (anonymous):

why you changed 8 into a positive #

OpenStudy (anonymous):

That is because I added 8 to both sides. x/7 - 8 = -7 x/7 - 8 + 8 = -7 + 8 x/7 = 1 7x/7 = 7 x = 7

OpenStudy (anonymous):

which 2 sides?

OpenStudy (anonymous):

Both sides of the equal to (=) sign. It is like a measuring scale. When you add, subtract, multiply and divide quantities on either side of the "=" sign, you keep the equation still in balance.

OpenStudy (anonymous):

that makes sense but what 2 side did you add 8 to?

OpenStudy (anonymous):

x/7 - 8 is considered one side. -7 is considered the other side.

OpenStudy (anonymous):

x/7 - 8 = -7 x/7 - 8 + 8 = -7 + 8 x/7 = 1 7x/7 = 7 x = 7

OpenStudy (anonymous):

so you solve the first side and add it by 8 and then you add -7 by 8?

OpenStudy (anonymous):

You are manipulating both sides at the same time. The purpose is to retain just "x" on one side and get all constants on to the other side. That is how you can find the value of "x", which is what the problem asked you to do (solve for x => find the value of x).

OpenStudy (anonymous):

can you explain that easier? i don't understand what you are saying?

OpenStudy (anonymous):

I am sorry, I don't know what grade you are in and who is teaching you this stuff at school. But, I can't help further.

OpenStudy (anonymous):

i am in 8th grade

OpenStudy (anonymous):

thanks 4 trying, but i have 2 eat soon. will you be on later?